Wednesday 02/15/12

9:45 A — 11:00P

  • Etching Dm0-1 slides (25-30 min) for STORM.  (Slide 69 in records)
  • Flip fly cage, 10A
  • Washing out primary antibody in PcG stained embryos
  • in situs for PcG stained embryos: hox and en locus (see records)
  • organized and sent NF results on PcG mitosis (see post)
  • forgot to collect virgins this morning.
  • send fedex info to Dorsett lab for Rad21 antibody
  • start application for SDB membership
  • STORM imaging slide H1 Dm01 (#69)  2.9 hr run
  • Meeting with DNA FISH collaborators (Ting Wu lab).
  • STORM imaging slide H2: (#70).  12-hr run O/N, may get stopped at hour 10 for GD’s time.
  • Moved probe library into -20C downstairs. Kept most used probes upstairs.
  • clear and collect phenotypic virgins (to remake snail BAC rescue).
This entry was posted in Summaries and tagged , , , . Bookmark the permalink.